Show that z(sqrt-2) is non ufd
1
Expert's answer
2020-04-13T16:49:36-0400

The given ring is "R=\Z[\sqrt{-2}]" .

Your question is wrong , Since we known that "R" is a Euclidean domain by defining a norm on "R" ,i,e,"N(a+b\sqrt{-2})=a^2+2b^2" and every Euclidean domain is PID .

Again every PID is UFD.

Hence "R" is a UFD.